Homer (age 68) and his wife Jean (age 70) file a joint return. They furnish all of the support of Luther (Homer's 90-year-old father), who lives with them. For 2011, they received $6,000 of interest income on city of Chicago bonds and interest income on corporate bonds of $48,000. Compute Homer’s and Jean's taxable income for 2011.
1
Expert's answer
2014-06-18T09:31:10-0400
Taxable income = adjusted income - (deductions + allowance for exemptions)

Interest income was $6,000.
Interest from corporate bonds was $48,000.
Adjusted income = 6,000 + 48,000 = 54,000.
In 2011 standard deduction was $11,600 for married filing jointly.
Additional standard deduction in 2011 based on age (65 or older) was $1,150 for married filing jointly per person, $2,300 for Homer and Jean. For father the additional standard deduction will be $1,450 as for single. Total amount will be $3,750.
TI = 54,000 - (11,600 + 3,750) = 38,650.